OpenStudy (anonymous):

if x,4,y are successive terms in an arithmetic sequence and x,3,y are successive terms in a geometric sequence, calculate 1/x + 1/y.

OpenStudy (shubhamsrg):

1/x + 1/y = (x+y)/(xy) now in AP ,, (x+y)/2 =4 and in GP ,, sqrt(xy) = 3 just substitute values..

OpenStudy (anonymous):

Can you continue? I still am n't too sure

OpenStudy (shubhamsrg):

from my 1st eqn,, x+y = 4*2 =8 from my 2nd eqn,, xy = 3^2 =9 so (x+y)/xy = 8/9 <= your ans

OpenStudy (anonymous):

ty.

OpenStudy (anonymous):

Though, how did you get x+y and sqrt(xy) in the first place

OpenStudy (shubhamsrg):

its arthimetic annd geometric mean if a,b,c are in AP,,then b=(a+c)/2 and if they are in GP,,then b = sqrt(ac) these are some simple formullaes which can be proved easily.. note general terms of AP are a,a+d,a+2d ... and of GP are a,ar, ar^2 ,ar^3 ...

OpenStudy (anonymous):

I see. Thanks again.

OpenStudy (anonymous):

Wait. Sorry to bother again. But how do you know x+y = 1/x and sqrt(xy) = 1/y?

OpenStudy (shubhamsrg):

nonono.. 1/x and 1/y are not equal to x+y and sqrt(xy) respectively,, but 1/x + 1/y = (x+y)/(xy) just take LCM on LHS and do the simple maths..

OpenStudy (anonymous):

Ok. Got it. I wasn't thinking hard enough >.<